OpenStudy (anonymous):

simplify the radical expression. Please help A. Square root of 25x^2 B. Square root of 125h^4

OpenStudy (anonymous):

sqrt(25x^2) = sqrt(25) * sqrt(x^2) = 5sqrt(x^2)

OpenStudy (anonymous):

the square root of twenty five is five, and the square root of x squared is x so the square root of twenty five x squared is five x

OpenStudy (anonymous):

more succinctly written as \[\sqrt{25x^2}=5x\]

OpenStudy (anonymous):

\[\sqrt{h^4}=h^2\] and \(\sqrt{125}=\sqrt{25\times 5}=\sqrt{25}\times \sqrt5=5\sqrt5\)making \[\sqrt{125h^4}=5h^2\sqrt5\]

OpenStudy (anonymous):

ah i think i understand thanks so much! @satellite73